OpenStudy (anonymous):

1.6x – 3.38 = 5.81x + 4.67 what is x?

OpenStudy (anonymous):

put x on one side constants on the other and then solve it

OpenStudy (anonymous):

5.81x-1.6x=-3.38-4.67 4.21x=-8.05 x=-1.912114014